### v1

@@ 4.3 The Class of ($s,S$) Policies
It remains to bound $\mathbb{E}_{\mathbf{d}_{1},\ldots,\mathbf{d}_{N}\sim\mathcal{D}}\left[\mathrm{GE}\left({\cal L}(\Pi_{(s,S)})\right)\right]$ from below. Note that for binary function classes, the VC-dimension provides both lower and upper bounds on the expected generalization error (e.g., Theorem 6.8 of Shalev-Shwartz and Ben-David 2014). In other words, the learnability of a binary-valued function class is equivalent to the finiteness of its VC-dimension. By contrast, a lower bound on the Pseudo-dimension of a function classes does not necessarily translate to a lower bound on the corresponding generalization error. In fact, the learnability of real-valued function classes is established through the finiteness of Pseudoγ-dimension, rather than Pseudo-dimension, and the Pseudoγ-dimension can provide both lower and upper bounds on the expected generalization error in certain contexts (e.g., Kearns and Schapire 1994, Alon et al. 1997). Therefore, we have to analyze a lower bound on the more complex Pseudoγ-dimension of ${\cal L}(\Pi_{(s,S)})$. Our next result presents such a lower bound on the Pseudoγ-dimension for some constant $\gamma>0$ by constructing an instance. Our proof involves the celebrated Prime Number Theorem and the asymptotic property of the first Chebyshev function.
@@ 4.4 The Class of Non-Stationary Base-Stock Policies
Of course, the simple argument above only works for the perishable setting. To establish an upper bound of $O(\sqrt{1/N})$ for the expected generalization error of $(S^{t})$ policies in the original, non-perishable setting, we now analyze the stronger notion of Pseudoγ-dimension. We use $y^{t}\left((S^{t^{\prime}})_{t^{\prime}=1}^{T+L},\mathbf{d}\right)$ to denote the inventory level after replenishment but before seeing demand in period $t$ under policy $(S^{t^{\prime}})_{t^{\prime}=1}^{T+L}$ on a given demand sequence $\mathbf{d}$. Note that $y^{t}\left((S^{t^{\prime}})_{t^{\prime}=1}^{T+L},\mathbf{d}\right)=y^{t}\left((S^{t^{\prime}})_{t^{\prime}=1}^{t-L},\mathbf{d}\right)\in[-LU,H]\subseteq[-(L+1)U,(L+1)U]$ for any $(S^{t^{\prime}})_{t^{\prime}=1}^{T+L}\in\Pi_{(S^{t})}$ and $\mathbf{d}\in[0,U]^{T+L}$. We further define a class of functions representing the normalized inventory level for $t=L+1,\ldots,T+L$:
$\mathcal{Y}^{t}(\Pi_{(S^{t})})\triangleq$$\left\{y^{t}\left((S^{t^{\prime}})_{t^{\prime}=1}^{t-L},\mathbf{d}\right)/((L+1)U):(S^{t^{\prime}})_{t^{\prime}=1}^{t-L}\in[0,H]^{t-L}\right\}.$
In other words, the function class $\mathcal{Y}^{t}(\Pi_{S^{t}})$ takes values in the interval $[-1,1]$.
@@ 6 Conclusion
This paper adopts a novel approach by analyzing generalization guarantees of several well-known classes of inventory policies, leveraging the celebrated VC theory. Both our theory and simulations reveal important managerial insights into the principle of ?learning less is more? in inventory management. Depending on the amount of data available, it may be beneficial to restrict oneself to a simpler, albeit suboptimal, class of inventory policies. Moreover, the number of parameters in a policy class may not accurately measure overfitting. For example, the generalization error for $(S^{t})$ policies is comparable to that of $(s,S)$ policies, despite having significantly more parameters.
Our work has the potential for broader, significant impacts. In recent years, many studies have applied state-of-the-art deep learning approaches to solve large-scale inventory problems. We believe our structured approach of learning base-stock levels instead of order quantities can potentially enhance existing deep learning techniques, which is particularly valuable in retailing where relevant data is often limited. In other words, it is important to leverage the domain knowledge from 70 years of inventory theory. The idea of leveraging academic knowledge and restricting to policy structures that have been well-established by theory can potentially translate to other problems too.
@@
The authors thank Matias Alvo for providing useful comments on an early draft.

### v2

@@ 4.3 The Class of ($s,S$) Policies
It remains to bound $\mathbb{E}_{\mathbf{d}_{1},\ldots,\mathbf{d}_{N}\sim\mathcal{D}}\left[\mathrm{GE}\left({\cal L}(\Pi_{(s,S)})\right)\right]$ from below. Note that for binary function classes, the VC-dimension provides both lower and upper bounds on the expected generalization error (e.g., Theorem 6.8 of Shalev-Shwartz and Ben-David 2014). In other words, the learnability of a binary-valued function class is equivalent to the finiteness of its VC-dimension. By contrast, a lower bound on the Pseudo-dimension of real-valued function classes does not necessarily translate to a lower bound on the corresponding generalization error. In fact, the learnability of real-valued function classes is established through the finiteness of Pseudoγ-dimension, rather than Pseudo-dimension, and the Pseudoγ-dimension can provide both lower and upper bounds on the expected generalization error in certain contexts (e.g., Kearns and Schapire 1994, Alon et al. 1997). Therefore, we have to analyze a lower bound on the more complex Pseudoγ-dimension of ${\cal L}(\Pi_{(s,S)})$. Our next result presents such a lower bound on the Pseudoγ-dimension for some constant $\gamma>0$ by constructing an instance. Our proof involves the celebrated Prime Number Theorem and the asymptotic property of the first Chebyshev function.
@@ 4.4 The Class of Non-Stationary Base-Stock Policies
Of course, the simple argument above only works for the perishable setting. To establish an upper bound of $O(\sqrt{1/N})$ for the expected generalization error of $(S^{t})$ policies in the original, non-perishable setting, we now analyze the stronger notion of Pseudoγ-dimension. We use $y^{t}\left((S^{t^{\prime}})_{t^{\prime}=1}^{T+L},\mathbf{d}\right)$ to denote the inventory level after replenishment but before seeing demand in period $t$ under policy $(S^{t^{\prime}})_{t^{\prime}=1}^{T+L}$ on a given demand sequence $\mathbf{d}$. Note that $y^{t}\left((S^{t^{\prime}})_{t^{\prime}=1}^{T+L},\mathbf{d}\right)=y^{t}\left((S^{t^{\prime}})_{t^{\prime}=1}^{t-L},\mathbf{d}\right)\in[-LU,(L+1)U]\subseteq[-(L+1)U,(L+1)U]$ for any $(S^{t^{\prime}})_{t^{\prime}=1}^{T+L}\in\Pi_{(S^{t})}$ and $\mathbf{d}\in[0,U]^{T+L}$. We further define a class of functions representing the normalized inventory level for $t=L+1,\ldots,T+L$:
$\mathcal{Y}^{t}(\Pi_{(S^{t})})\triangleq$$\left\{\frac{y^{t}\left((S^{t^{\prime}})_{t^{\prime}=1}^{t-L},\mathbf{d}\right)}{(L+1)U}:(S^{t^{\prime}})_{t^{\prime}=1}^{t-L}\in[0,H]^{t-L}\right\}.$
Note that the function class $\mathcal{Y}^{t}(\Pi_{S^{t}})$ takes values in the interval $[-1,1]$.
@@ 6 Conclusion
This paper analyzes generalization guarantees of several well-known classes of inventory policies on correlated demand sequences, leveraging the celebrated VC theory. Our derivation of learning-theoretic dimensions for inventory trajectories diverges from traditional data-driven inventory literature, and our problem-specific analysis for inventory also appears new in the context of data-driven algorithm design. Both our theory and simulations reveal important managerial insights into the principle of ?learning less is more? in inventory management: depending on the amount of data available, it may be beneficial to restrict oneself to a simpler, albeit suboptimal, class of inventory policies. Moreover, we find surprisingly that $(s,S)$ policies are harder to learn than $(S^{t})$ policies, despite having 2 instead of $T$ parameters. Our bound for learning $(S^{t})$ policies based on the powerful VC theory is better than some existing bounds for $(S^{t})$ policies even if they assume independent demands over time.
All in all, our work advocates for learning well-defined policy classes, which has the potential for broader, significant impacts. Indeed, many recent studies have applied state-of-the-art deep learning approaches to solve large-scale inventory problems. We believe our structured approach of learning base-stock levels instead of order quantities can potentially enhance existing deep learning techniques, which is particularly valuable in retailing where relevant data is often limited. In other words, it is important to leverage the domain knowledge from 70 years of inventory theory. The idea of leveraging academic knowledge and restricting to policy structures that have been well-established by theory can potentially translate to other problems too.
@@
The authors thank Matias Alvo, Baris Ata, John R. Birge, René Caldentey, Xin Chen, Yilun Chen, David A. Goldberg, Simai He, Zhiyi Huang, Chengyi Lyu, Hanzhang Qin, Cong Shi, Ellen Vitercik, Amy R. Ward, and Sean Zhou for insightful discussions and/or constructive comments on an early draft.